# Alleviating Non-identifiability: a High-fidelity Calibration Objective for Financial Market Simulation with Multivariate Time Series Data

### v3

@@ 3.2 Alleviating the Non-identifiability with Multiple Features
Suppose the observed time series generated by the real social system is multivariate containing $K$ features, i.e., $\hat{\mathbf{X}}_{T}=\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$, which is true for the financial markets. And the simulation model $M(\omega)$ can also generate a multivariate simulated data $\mathbf{X}_{T}=\{\mathbf{X}_{T}^{k}\}_{k=1}^{K}$ with the same $K$ features. Given that existing discrepancy metrics $D$ only computes for two univariate time series, there are in total $K$ individual calibration tasks: $\min_{w}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),k=1,\cdots,K.$
Note that, given any two parameters $\omega_{1},\omega_{2}\in\Omega^{D,k}$, only they are simultaneously non-identifiable in all $K$ individual calibration tasks, i.e., $\omega_{1},\omega_{2}\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$, are they non-identifiable in the general calibration task with respect to multivariate $\hat{\mathbf{X}}_{T}$. Otherwise, their discrepancy to $\hat{\mathbf{X}}_{T}$ can be distinguished in at least one feature and thus are identifiable. The search space $\Omega^{D,k}$ is identical across all calibration tasks and is uniformly represented as $\Omega^{D}$. That is to say, only the intersection of the $K$ individual non-identifiable set $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$ can be regarded as the non-identifiable set in the multivariate time series setting, which is defined as
@@
Let $P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)$ be the non-identifiability to the multivariate time series with $K$ features. It it defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D}$ while it falls into all $K$ individual non-identifiable set $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$. And we have
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)=\frac{\left|\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D}\right|}.$ (4)
Since $\frac{\left|\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D}\right|}\leq\min_{k}\frac{\left|\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D,k}\right|}$, we have $P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)\leq\min_{k}P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right)$. This equality holds only if the $K$ non-identifiable sets are fully overlapped. This can be caused that the $K$ features are fully dependent that the time series data of one feature can be deduced from other features. This suggests that the selected features should be as diverse as possible. Therefore, with proper choice of the features and $D$, the non-identifiability of calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$ is smaller than that of the univariate $\hat{\mathbf{X}}_{T}^{k}$. In other words, calibrating multiple features can alleviate the non-identifiability issue. Next, we estimate how fast it decreases with $K$ as follows.
@@
$=\frac{\left|\mathcal{S}^{D,1,\epsilon}\right|}{\left|\Omega^{D}\right|}\cdot\frac{\left|\cap_{i=1}^{2}\mathcal{S}^{D,i,\epsilon}\right|}{\left|\mathcal{S}^{D,1,\epsilon}\right|}\cdots\frac{\left|\cap_{i=1}^{K}\mathcal{S}^{D,i,\epsilon}\right|}{\left|\cap_{i=1}^{K-1}\mathcal{S}^{D,i,\epsilon}\right|}$
[…]
@@
[…]
3.3 Aggregating Multiple Features via Maximization
Note that, the above discussions require the multivariate time series $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$ being calibrated separately as $K$ univariate time series with existing discrepancy metrics $D$. This results in $K$ individual calibration tasks of $D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),k=1,\cdots,K$. How to jointly calibrate the parameter to satisfy the $K$ individual tasks simultaneously? This section proposes a new objective function based on the $K$ individual objectives, so that the practical alleviation of the non-identifiability follows the above theory.
Let $F$ be the proposed new objective function used for calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$, and let $\mathcal{S}^{F,\epsilon}$ denote the associated non-identifiable set defined by this objective function. To achieve $\mathcal{S}^{F,\epsilon}=\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$, the proposed way is to ensure every location in $S^{F}$ to be the worst value among all $K$ non-identifiable sets $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$ (see Figure 3 for illustration). In this case, any parameter $\omega$ that is not simultaneously in all $K$ non-identifiable sets will make it fall outside of the $\mathcal{S}^{F,\epsilon}$. This is because that the parameters outside the non-identifiable set is even worse than those who are non-identifiable from the ground-truth parameter $\omega^{*}$.
Figure 3: Optimizing the maximization of two functions equals to searching in their intersections.
@@ 3.3 Aggregating Multiple Features via Maximization
Theorem 2 (The utility of new objective function $F$)
Let $F$ aggregate the $K$ individual calibration tasks via maximization, i.e.
$F(\hat{\mathbf{X}}_{T},M(w)=\mathbf{X}_{T})$ (6)
$=$ $\max_{k\in\{1,\cdots,K\}}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),$
then minimizing $F$ achieves $\mathcal{S}^{F,\epsilon}=\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$.
The proof of Theorem 2 is provided in appendix.
4 Empirical Studies
@@
4.1 Experimental Settings
To conduct a calibration process, 4 terms should be specified according to Eq. (6), i.e., the discrepancy metric $D$, the simulation model $M(\omega)$, the optimization algorithm $\rm{min}_{\omega}$, and the $K$ features $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$.
The discrepancy metric. The Wasserstein distance is employed as $D$ since it provides geometric interpretability, and tail sensitivity  [6, 5]. Given two univariate time series $\mathbf{X}_{T}^{k}$ and $\hat{\mathbf{X}}_{T}^{k}$, the Wasserstein distance is defined as
@@ 4.1 Experimental Settings
$=$ $\int_{-\infty}^{+\infty}\left|\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\hat{\mathbf{X}}_{t}^{k}\leq x)-\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\mathbf{X}_{t}^{k}\leq x)\right|\,\mathrm{d}x,$
@@
where $\mathcal{I}(\cdot)$ is an indicator function that returns 1 if the input event is true, otherwise it returns 0.
@@
The simulation model. This work considers the Preis-Golke-Paul-Schneid (PGPS) model as the simulation model, due to its popularity in recent financial market simulation works  [25, 11]. The PGPS model simulates the market dynamics through interactions between two types of agents, i.e., 125 liquidity providers and 125 liquidity takers. The PGPS model has 6 hyper-parameters shared by all agents that need to be calibrated, i.e., $\omega=[\delta,\lambda_{0},C_{\lambda},\Delta_{s},\alpha,\mu]$. The detailed descriptions of the behaviors of the agents and how the hyper-parameters control them can be found in Appendix.
@@
The optimization method. Since the order matchmaking process needs to align with the real market, the PGPS model should be run on the executable trading programs. As a result, the calibration objective function is non-differentiable and the calibration process often has to resort to non-derivative optimization methods. In such contexts, both Bayesian optimization methods and evolutionary algorithms are particularly suitable. This work employs the well-known Particle Swarm Optimization (PSO) algorithm  [15] to optimize $\omega$ with respect to Eq. (6). The setting of PSO is in Appendix.
@@
  • $f_{5}$, the total volume of the best bid price at each $t$-th step;
  • $f_{6}$, the total volume of the best ask price at each $t$-th step.
By incorporating the 6 selected features, 6 objective functions are constructed based on Eq. (6), denoted as $F_{i},i=1,2,...,6$. Each $F_{i}$ indicates that the first $i$ features are used for calibration, i.e., $F_{i}=\max_{k\in\{1,\cdots,i\}}D(\hat{\mathbf{X}}_{T}^{k},M(\omega))$. In the following empirical studies, it is shown how the increasing number of features influences the simulation fidelity.
The test protocol. The test data contains 11 targeted time series to be calibrated, including 10 synthetic data and 1 real data. The 10 synthetic time series are generated by setting 10 different parameters of the PGPS model as shown in Appendix, each of which consists of 3600 time steps at the frequency of 1 second. The real data is chosen as 000001.sz from Shenzhen Stock Exchange of China, consisting of 1200 time steps at a frequency of 3 seconds from 9:30 a.m. to 10:30 a.m. of a day in 2019. The time budget of $10^{4}$ simulation evaluations is allowed for the calibration of each targeted data. That is, the PSO runs for 250 iterations with population size 40, and the best parameter with the lowest objective function value is applied into PGPS model to simulate data as $\mathbf{X}_{T}$. The performance of the calibration is quantitatively validated using the Wasserstein (W) distance and the Mean Square Error (MSE) commonly seen in time series analysis. Both metrics are applied to measure the averaged discrepancy between the synthetic data and simulated data on 6 features, i.e., $\frac{\sum_{k=1}^{6}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k})}{6}$, where $D$ here indicates W or MSE. Each feature is normalized with the min-max method, respectively.
4.2 Results and Analysis
@@
Alleviation of non-identifiability. Figure 4 illustrates the changes of the non-identifiable sets in the 2-dimensional parameter space of $[\alpha,\mu]$. This is done by three steps: 1) a target data of 3600 time steps is generated using the PGPS model with the recommended default parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$  [11]; 2) On this default setting, 10000 combinations are created by sampling in $[\alpha,\mu]$ in grid while keeping the values of other 4 parameters unchanged; 3) By using $F$ to measure the descrepancy between the target data and the simulated data of each of the 10000 combinations, it obtains the objective values of the 10000 combinations on each $F_{i}$, which are depicted as Figure 4. The 6 figures in Figure 4 are depicted according to the objective values of $F_{1}$, $F_{2}$,…,$F_{6}$, respectively. Only the combinations whose objective values smaller than $\epsilon=0.1$ are depicted in grey, otherwise they will be depicted in blue. For example, in the third figure marked with $F_{3}$, the combinations depicted in grey are those whose discrepancy values smaller than $\epsilon=0.1$ on all the first 3 features $\{f_{i}\}_{i=1}^{3}$. By this means, the grey points indicate the non-identifiable set as they are indistinguishable with $\omega^{*}$ (marked as the red dot). As can be seen that, with the increase of the features, the grey non-identifiable set, decreases significantly. When 4 are incorporated in $F$, there exists only the optimum, indicating that the non-identifiability issue is indeed alleviated very effectively.
[…]

### v7

@@ III-B Alleviating the Non-identifiability with Multiple Features
Suppose the observed time series generated by the real social system is multivariate containing $K$ features, i.e., $\hat{\mathbf{X}}_{T}=\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$, which is true for the financial markets. And the simulation model $M(\omega)$ can also generate a multivariate simulated data $\mathbf{X}_{T}=\{\mathbf{X}_{T}^{k}\}_{k=1}^{K}$ with the same $K$ features. Given that existing discrepancy metrics $D$ only compute for two univariate time series, there are naturally in total $K$ individual calibration tasks: $\min_{w}D(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}_{T}^{k}),k=1,\cdots,K.$
Note that, given any two parameters $\omega_{1},\omega_{2}\in\Omega^{D,k}$, only they are simultaneously non-identifiable in all $K$ individual calibration tasks, i.e., $\omega_{1},\omega_{2}\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$, are they non-identifiable in the general calibration task with respect to multivariate $\hat{\mathbf{X}}_{T}$. Otherwise, their discrepancy to $\hat{\mathbf{X}}_{T}$ can be distinguished in at least one feature and thus are identifiable. The search space $\Omega^{D,k}$ is identical across all calibration tasks and is uniformly represented as $\Omega^{D}$. That is to say, only the intersection of the $K$ individual non-identifiable set $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$ can be regarded as the non-identifiable set in the multivariate time series setting, which is defined as follows:
@@
Let $P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)$ be the non-identifiability to the multivariate time series with $K$ features. It is defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D}$ while it falls into all $K$ individual non-identifiable set $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$. And we have
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)=\frac{\mu\left(\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\right)}{\mu\left(\Omega^{D}\right)}.$ (4)
Since
$\frac{\mu\left(\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\right)}{\mu\left(\Omega^{D}\right)}\leq\min_{k}\frac{\mu\left(\mathcal{S}^{D,k,\epsilon}\right)}{\mu\left(\Omega^{D,k}\right)}$ (5)
we have
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)\leq\min_{k}P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right).$ (6)
This equality holds only if the $K$ non-identifiable sets are fully overlapped. This can be caused by the fact that the $K$ features are fully dependent and that time series data of one feature can be deduced from other features. This suggests that the selected features should be as diverse as possible. Therefore, with proper choice of the features and $D$, the non-identifiability of calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$ is smaller than that of the univariate $\hat{\mathbf{X}}_{T}^{k}$. In other words, calibrating multiple features can alleviate the non-identifiability issue. Next, we estimate how fast it decreases with $K$ as follows.
@@
$=\frac{\mu\left(\mathcal{S}^{D,1,\epsilon}\right)}{\mu\left(\Omega^{D}\right)}\cdot\frac{\mu\left(\bigcap_{i=1}^{2}\mathcal{S}^{D,i,\epsilon}\right)}{\mu\left(\mathcal{S}^{D,1,\epsilon}\right)}\cdots\frac{\mu\left(\bigcap_{i=1}^{K}\mathcal{S}^{D,i,\epsilon}\right)}{\mu\left(\bigcap_{i=1}^{K-1}\mathcal{S}^{D,i,\epsilon}\right)}$ (7a)
$=\beta_{1}\cdot\beta_{2}\cdot\beta_{3}\cdots\beta_{K}$ (7b)
$\leq\left(\frac{\sum_{i=1}^{K}\beta_{i}^{2}}{K}\right)^{\frac{K}{2}}.$ (7c)
The proof of (7c) is provided in appendix A-A. In (7b), $\beta_{1}$ represents the non-identifiability of feature 1, i.e., the probability that any $\omega\in\mathcal{S}^{D,1,\epsilon}$ falls into $\Omega^{D}$. $\beta_{i},i=2,...,K$ denotes the probability that any parameter $\omega\in\cap_{k=1}^{i}\mathcal{S}^{D,k,\epsilon}$ falls into $\cap_{k=1}^{i-1}\mathcal{S}^{D,k,\epsilon}$. In other words, $\beta_{i}$ means the overlapping ratio between $\mathcal{S}^{D,i,\epsilon}$ and $\cap_{k=1}^{i-1}\mathcal{S}^{D,k,\epsilon}$, and $\beta_{i}\leq 1$. Especially, $\beta_{i}=1$ only if $\cap_{k=1}^{i-1}\mathcal{S}^{D,k,\epsilon}\subseteq\mathcal{S}^{D,i,\epsilon}$, indicating that the selected $i$-th feature is fully dependent on the first $i-1$ features.
The order of features in (7b) can be arbitrary, but it influences the upper bound of the multivariate non-identifiability, which is defined as (7c). That is,
@@
This is intuitive that if $\beta_{K}^{2}<\frac{{\sum}_{i=1}^{K-1}\beta_{i}^{2}}{K-1}$, the base of (7c) will not increase. This theorem not only provides an exponential decreasing upper bound for the multivariate non-identifiability but theoretically suggests a rule to select the features. The proof of Theorem 1 is in appendix A-B.
III-C Aggregating Multiple Features via Maximization
Note that the above discussions require the multivariate time series $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$ being calibrated separately as $K$ univariate time series with existing discrepancy metrics $D$. This results in $K$ individual calibration tasks of $D(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}_{T}^{k}),k=1,\cdots,K$. How to jointly calibrate the parameter to satisfy the $K$ individual tasks simultaneously? This section proposes a new objective function based on the $K$ individual objectives so that the practical alleviation of the non-identifiability follows the above theory.
Let $F$ be the proposed new objective function used for calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$, and let $\mathcal{S}^{F,\epsilon}$ denote the associated non-identifiable set defined by this objective function. To achieve $\mathcal{S}^{F,\epsilon}=\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$, the proposed way is to ensure every location in $S^{F}$ to be the worst value among all $K$ non-identifiable sets $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$ (see Fig. 4 for illustration). In this case, any parameter $\omega$ that is not simultaneously in all $K$ non-identifiable sets will make it fall outside of the $\mathcal{S}^{F,\epsilon}$. This is because the parameters outside the non-identifiable set are even worse than those that are non-identifiable from the ground-truth parameter $\omega^{*}$.
@@ III-C Aggregating Multiple Features via Maximization
Theorem 2
(The utility and uniqueness of new objective function $F$) Let $F$ aggregate the $K$ individual calibration tasks via maximization, i.e.
$F\left(\hat{\mathbf{X}}_{T},M(\omega)=\mathbf{X}_{T}\right)$ (8)
$=$ $\max_{k\in\{1,\cdots,K\}}D\left(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}_{T}^{k}\right),$
then minimizing $F$ uniquely achieves $\mathcal{S}^{F,\epsilon}=\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$.
The proof of Theorem 2 is provided in appendix A-C.
Figure 4: Optimizing the maximization of two functions equals searching in their intersections.
IV Empirical Studies
@@
TABLE I: The ranges for randomly sampling parameters $\omega$ to generate the synthetic data
Para. Range Remarks
$\delta$ $[0.00,0.050]$ The probability for each liquidity taker to cancel an untraded order.
$\lambda_{0}$ $[50.00,300.000]$ Controlling the price of each limited order.
$C_{\lambda}$ $[1.00,50.000]$ Controlling the price of each limited order.
$\Delta_{S}$ $[0.00,0.005]$ Controlling the probability of the side of a market order (ask or bid).
$\alpha$ $[0.05,0.500]$ The probability of each liquidity provider submitting a limited order.
$\mu$ $[0.00,0.050]$ The probability of each liquidity taker submitting a market order.
IV-A Experimental Settings
To conduct a calibration process, 4 terms should be specified according to (8), i.e., the discrepancy metric $D$, the simulation model $M(\omega)$, the optimization algorithm $\rm{min}_{\omega}$, and the $K$ features $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$.
The discrepancy metric. The Wasserstein distance is employed as $D$ since it provides geometric interpretability and tail sensitivity [29, 30]. Given two univariate time series $\mathbf{X}_{T}^{k}$ and $\hat{\mathbf{X}}_{T}^{k}$, the Wasserstein distance is defined as
@@ IV-A Experimental Settings
$=$ $\int_{-\infty}^{+\infty}\left|\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\hat{\mathbf{X}}_{t}^{k}\leq x)-\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\mathbf{X}_{t}^{k}\leq x)\right|\rm{d}\textit{x},$
@@
where $\mathcal{I}(\cdot)$ is an indicator function that returns 1 if the input event is true. Otherwise, it returns 0.
@@
The simulation model. This work considers the Preis-Golke-Paul-Schneid (PGPS) model as the simulation model due to its popularity in recent FMS works [13, 8]. The PGPS model simulates the market dynamics through interactions between two types of agents, i.e., 125 liquidity providers and 125 liquidity takers. The PGPS model has 6 hyper-parameters shared by all agents that need to be calibrated, i.e., $\omega=[\delta,\lambda_{0},C_{\lambda},\Delta_{s},\alpha,\mu]$. At the $t$-th time-step, each liquidity provider submits a limited order at a fixed probability $\alpha$ with the default volume equals 1. The probability of a limited order being bid side or ask side equals 0.5. Each liquidity taker submits a market order at a fixed probability $\mu$ with the default volume equals 1. The probability of a market order being bid side or ask side is $q_{taker}(t)$ or $1-q_{taker}(t)$, respectively. The probability $q_{taker}(t)$ is specified by a mean-reverting random walk with mean equals 0.5, mean reversion probability equals $0.5+|q_{taker}(t)-0.5|$, and the increment sizes towards mean equals ±$\pm\Delta_{s}$. Moreover, the liquidity taker has a probability $\delta$ to cancel its untraded limited order. Let $p_{a}(t)$ and $p_{b}(t)$ represent the best ask and bid price, respectively. The price of a market order is determined by the market, i.e., the best price at the opposite side of the LOB. The price of a limit order is determined as $p=p_{s}(t)+\lambda(t)\log(u)+s$ where $p_{s}(t)=p_{a}(t)$, $s=1$ for an ask order, and $p_{s}(t)=p_{b}(t)$, $s=-1$ otherwise, with $u\sim U(0,1)$. Here $\lambda(t)$ is a time-variant order placement depth parameter and is calculated as
$\lambda(t)=\lambda_{0}\left(1+\frac{\left|q_{taker}(t)-\frac{1}{2}\right|}{\sqrt{<q_{taker}(t)-\frac{1}{2}>^{2}}}C_{\lambda}\right)$ (10)
$\sqrt{<q_{taker}(t)-\frac{1}{2}>^{2}}$ is a pre-computed value. Before initiating the simulation, it is computed in advance through $10^{5}$ Monte Carlo iterations, ensuring its convergence to the accurate standard deviation.
@@
The optimization method. Since the order matchmaking process needs to align with the real market, the PGPS model should be run on the executable trading programs. As a result, the calibration objective function is non-differentiable and the calibration process often has to resort to non-derivative optimization methods. In such contexts, both Bayesian optimization methods and evolutionary algorithms are particularly suitable. This work employs the well-known Particle Swarm Optimization (PSO) algorithm [31], which has been frequently adopted to train the financial simulation or prediction models[6, 21, 32], to optimize $\omega$ with respect to (8). For simplicity, the standard version of PSO [33] is considered. The pseudocode for calibration with PSO is shown in Algorithm 1.
Algorithm 1 Calibration with PSO
Input: Number of Iterations to run $\kappa$
Parameter: The hyper-parameters of PSO
Output: The optimal candidate $\omega^{*}$
1:  Let $t=0$.
2:  Initialize a population of $N$ particles $\{\omega_{i}\}^{N}_{i=1}$.
3:  Simulate each of $N$ particles with $M(\omega)$ to obtain $N$ simulated data.
4:  Calculate the performance of each of $N$ particles as the Wasserstein distance.
5:  Set $\omega^{*}$ to the best performed candidate.
6:  while $t<\kappa$ do
7:     Generate a new population of $N$ particles $\{\omega_{i}\}^{N}_{i=1}$ using PSO operators.
8:     Simulate each of $N$ particles with $M(\omega)$ to obtain $N$ simulated data.
9:     Calculate the performance of each of $N$ particles as the Wasserstein distance.
10:     Update $\omega^{*}$ to the best performed candidate.
11:  end while
12:  return $\omega^{*}$
The PSO hyper-parameters, population size $N$, inertia weight $\omega$, cognitive/social coefficients $c_{1}$, $c_{2}$ are set to $N=40$, $\omega=0.8$, $c_{1}=c_{2}=0.5$, chosen from ranges $N\in[30,50]$, $\omega\in[0.6,1.0]$, $c_{1},c_{2}\in[0.3,0.7]$ based on experiments.
TABLE II: Parameter settings of generating 10 synthetic data with the PGPS model
Parameters $\mathbf{\delta}$ $\mathbf{\lambda_{0}}$ $\mathbf{C_{\lambda}}$ $\mathbf{\Delta_{S}}$ $\mathbf{\alpha}$ $\mathbf{\mu}$
data 1 0.025 100 10 0.001 0.15 0.025
data 2 0.002 200 10 0.002 0.1 0.03
data 3 0.05 152 45 0.003 0.13 0.05
data 4 0.02 288 27 0.003 0.11 0.04
data 5 0.04 62 35 0.003 0.12 0.04
data 6 0.01 171 15 0.0015 0.15 0.03
data 7 0.033 114 14 0.0033 0.1 0.047
data 8 0.01 129 30 0.0017 0.05 0.03
data 9 0.02 62 2 0.003 0.14 0.02
data 10 0.01 87 23 0.001 0.09 0.05
@@
  • $f_{5}$, the volume of the best bid price at each $t$-th step;
  • $f_{6}$, the volume of the best ask price at each $t$-th step.
By incorporating the 6 selected features, 6 objective functions are constructed based on (8), denoted as $F_{i},i=1,2,...,6$. Each $F_{i}$ indicates that the first $i$ features are used for calibration, i.e., $F_{i}=\max_{k\in\{1,\cdots,i\}}D(\hat{\mathbf{X}}_{T}^{k},M(\omega))$. The following empirical studies show how the increasing number of features influences the simulation fidelity.
The test protocol. The test data consists of 11 targeted time series to be calibrated, including 10 synthetic data sets and 1 real data set. To ensure sufficient diversity, the 10 synthetic time series are generated using the PGPS model by randomly sampling parameters within the ranges specified in table I, as suggested by [13]. The 10 sampled parameters are listed in table II. Each synthetic time series comprises 3600 time steps at a frequency of 1 second. The real data is chosen as 000001.sz from Shenzhen Stock Exchange of China, consisting of 1200 time steps at a frequency of 3 seconds from 9:30 a.m. to 10:30 a.m. of a day in 2019. The time budget of $10^{4}$ simulation evaluations is allowed for the calibration of each targeted data. That is, the PSO runs for 250 iterations with a population size of 40, and the best parameter with the lowest objective function value is applied to the PGPS model to simulate data as $\mathbf{X}_{T}$.
@@
Figure 7: The non-identifiability (NI) decreasing behaviors with the addition of new features.
Calibration performance is evaluated using the Wasserstein (W) distance and Mean Square Error (MSE), two common metrics in time series analysis. Both metrics measure the average discrepancy between synthetic and simulated data across 6 features, i.e., $\frac{\sum_{k=1}^{6}D(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}_{T}^{k})}{6}$, where $D$ denotes W or MSE. Each feature is normalized by the min-max method.
The computational platform. This work builds the simulation environment based on the Multi-Agent Exchange Environment (MAXE) simulator developed by the University of Oxford [35]. The experiments were run on a server with 500GB of memory, the Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz with 32 physical cores. For a single objective, PSO can be run in parallel across 40 simulators. The calibration of 6 objective functions for 10 synthetic data takes 48 to 60 hours in a data-parallel manner using the CPU.
IV-B Results on Alleviation of Non-identifiability
Fig. 5 illustrates the changes of the non-identifiable sets in the 2-dimensional parameter space of $[\alpha,\mu]$. This is done by three steps: 1) a target data of 3600 time steps is generated using the PGPS model with the recommended default parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$ [8]; 2) On this default setting, 10000 combinations are created by sampling in $[\alpha,\mu]$ in the grid while keeping the values of other 4 parameters unchanged; 3) By using $F$ to measure the discrepancy between the target data and the simulated data of each of the 10000 combinations, it obtains the objective values of the 10000 combinations on each $F_{i}$, which are depicted as Fig. 5. The 6 figures in Fig. 5 are depicted according to the objective values of $F_{1}$, $F_{2}$,…,$F_{6}$, respectively. Only the combinations whose objective values are smaller than $\epsilon=0.1$ are depicted in grey; otherwise, they will be depicted in blue. For example, in the third figure marked with $F_{3}$, the combinations depicted in grey are those whose discrepancy values are smaller than $\epsilon=0.1$ on all the first 3 features $\{f_{i}\}_{i=1}^{3}$. By this means, the grey points indicate the non-identifiable set as they are indistinguishable from $\omega^{*}$ (marked as the red dot). As can be seen, with the increase of the features, the grey non-identifiable set decreases significantly. When 4 are incorporated in $F$, only the optimum exists, indicating that the non-identifiability issue is indeed alleviated very effectively.
Fig. 6 further depicts how the non-identifiability decreases with the increasing numbers of features with different settings of $\epsilon$. This is done by calculating (7b) using the above 10000 combinations. Clearly, all the curves follow an exponentially decreasing manner. Besides, a larger value of $\epsilon$ is, the slower the corresponding curve decreases.